OpenStudy (anonymous):

pauls dad made a turkey pot pie for dinner on wednesday. the family ate 4/8 of the pie. on thursday after school, paul ate 2/16 of the pie for snack. what fraction of the pie remained

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

im bad at word problems

OpenStudy (anonymous):

forget the words. \[\frac{4}{8}\] is gone. \[\frac{2}{16}=\frac{1}{8}\] is gone, so how much is gone all together?

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

in other words, what is \[\frac{4}{8}+\frac{1}{8}\]?

OpenStudy (anonymous):

oh. one sec

OpenStudy (anonymous):

derp 5/8

OpenStudy (anonymous):

derp! lol

OpenStudy (anonymous):

ok so \[\frac{5}{8}\] is gone, what is left?

OpenStudy (anonymous):

probably easier if you think about this in english. five-eighths is gone, how many eights remain?

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

eiggggght?

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

hold the phone. you have 8 dollars and you spend 5 dollars, how much do you have left?

OpenStudy (anonymous):

3/8! I LITERALLY just figured it out

OpenStudy (anonymous):

great! (words not so bad once you get rid of them)

OpenStudy (anonymous):

lol